. Find the value of Z if the normal curve area between 0 and Z (positive) is 0.4332.
1
Expert's answer
2020-05-18T20:06:58-0400
"P(0<Z<z)=P(Z<z)-P(Z\leq0)="

"=P(Z<z)-0.5=0.4332"

"P(Z<z)=0.9332"

"z\approx1.50"